OpenStudy (exoexoexo):

the number of bacteria in a culture increases by 20% every 20 minutes. If there are 1,000 bacteria present at noon on a given day, to the nearest thousand, how many will be present at midnight of the same day?

OpenStudy (anonymous):

you can us \[p = p _{0}e ^{rt}\] P = Total population after time \[P_{o}\] = starting population r = % of growth rate T = Time in hours or years and e = Euler number

OpenStudy (exoexoexo):

what is an euler number?

OpenStudy (anonymous):

it's e on your calculator and it is a constant like pi e = 2.71828....

OpenStudy (anonymous):

alternatively you since this also creates a geometric sequence you can just calculate it like this \[1000(1.2)^{36}\]

OpenStudy (anonymous):

since noon to midnight is composed of 36, 20 minute intervals. and the common ratio is 1.2

OpenStudy (anonymous):

whichever way you like best :)

OpenStudy (danjs):

you see the form of this type of growth looks like Number Bacteria = (some start value at t=0) * (ratio of change) every 20 min, the bacteria are multiplied by that (ratio of change) , that is why this is an exponential growth Number Bacteria = (some start value at t=0) * (ratio of change)^(# of time intervals)

OpenStudy (danjs):

define noon to be time t=0, and then the bacteria = 1000 y = 1000 * (120/100)^x

OpenStudy (danjs):

The number increases to 120% of the previous, or, grows by 20%, every 20 min, the power x is then multiples of 20 min, or just say that is 1/3 hour, and let x be in hours Increases by 20% every 1/3 hour

OpenStudy (danjs):

so the power is then (3x) if you let x be hours, every hour, increases 3 times,